# Stability of $f(Q,B)$ gravity via Dynamical System Approach: a Comprehensive Bayesian Statistical Analysis

### v4

@@ Abstract
Abstract: In this work, we explore the cosmological stability of $f(Q,B)$ gravity using a dynamical system approach, where $Q$ denotes the nonmetricity scalar and $B$ represents the boundary term. We determine the model parameters of $f(Q,B)$ through Bayesian statistical analysis, employing Markov Chain Monte Carlo (MCMC) techniques. This analysis incorporates numerical solutions and observational data from Cosmic Chronometers (CC), the extended Pantheon+ dataset, and Baryonic Acoustic Oscillation (BAO) measurements. Our findings reveal a stable critical point within the dynamical system of the model, corresponding to the de Sitter phase, which is consistent with current observations of the Universe dominated by dark energy (DE) and undergoing late-time accelerated expansion. Additionally, we utilize Centre Manifold Theory to examine the stability of this critical point, providing deeper insights into the behavior of the model. The cosmological implications of $f(Q,B)$ gravity indicate a smooth transition in the deceleration parameters from deceleration to acceleration phase, underscoring the potential of the model to describe the evolution of the Universe. Our results suggest that the $f(Q,B)$ model presents a viable alternative to the standard $\Lambda$CDM model, effectively capturing the observed acceleration of the Universe and offering a robust framework for explaining the dynamics of cosmic expansion.
Keywords: Nonmetricity Scalar, Boundary Term, Cosmological Datasets, Dynamical System Analysis.
I Introduction
@@
The Lambda Cold Dark Matter ($\Lambda$CDM) model is widely regarded as the standard cosmological model, providing a fundamental framework for interpreting cosmic microwave background (CMB) data [1]. It incorporates cold dark matter, and late-time accelerated expansion driven by dark energy (DE) within a homogeneous and isotropic Universe. The cosmological constant $(\Lambda)$, integrated into Einstein’s field equations, has a robust theoretical basis. However, recent observations have highlighted discrepancies and tensions in $\Lambda$CDM model parameters across different cosmic epochs [2, 3], alongside persistent theoretical challenges associated with $\Lambda$ [4, 5]. These inconsistencies suggest that the $\Lambda$CDM model may not fully account for all missing physics. The discovery of the accelerating Universe expansion is one of the most surprising findings in cosmology, indicating either a breakdown of Einstein’s theory of gravity on cosmological scales or the dominance of DE with exotic properties. Cosmic acceleration has driven efforts to uncover its origin, with experiments measuring expansion and growth to one percent precision or better. General relativity (GR) is non-renormalizable at microscopic scales [6, 7], and the absence of a coherent quantum gravity theory necessitates modifications or extensions of GR [6, 8]. The standard approach involves developing gravity theories that encompass GR in specific scenarios while introducing additional degrees of freedom [9, 10, 11, 12, 13]. These theories arise from fundamental considerations of gravitational field structures and principles [14, 15, 16], with a focus on maintaining the principle of equivalence at quantum scales and addressing causal and geodesic structures. The metric requirement remains essential when considering gravity as a gauge theory.
@@
Novel classes of modified gravity theories have emerged, incorporating curvature, torsion, and nonmetricity scalars. These classes arise even though the unmodified theories are mathematically equivalent at the equation level. The key lies in the difference between the torsion scalar $T$ and the nonmetricity scalar $Q$, which deviates from the usual Levi-Civita Ricci scalar $\mathring{R}$ of GR due to additional terms: $\mathring{R}=-T+B$ and $\mathring{R}=Q+B$, respectively; where $B$ is the boundary term. All quantities denoted by $(\,\mathring{}\,)$ are computed in relation to the Levi-Civita connection $\mathring{\Gamma}$. The objects framed in GR can be identified by an over-circle symbol. A geometric trinity of gravity of second-order can be observed in $\mathring{R}$, $B-T$, $Q+B$, whereas $f(\mathring{R})$, $f(B-T)$, $f(Q+B)$ can be regarded as a geometric trinity of gravity of fourth-order [17, 18, 19]. Consequently, arbitrary functions $f(\mathring{R})$, $f(T)$, and $f(Q)$ no longer share a total derivative relationship. Furthermore, scalar fields can be introduced within this framework, leading to theories of scalar-tensor [20, 21], scalar torsion [22, 23, 24], and scalar nonmetricity [25, 26], each offering intriguing possibilities. Recently, Heisenberg [27] reviewed various cosmological models in $f(Q)$ gravity. Considering energy conditions, Banerjee et al. [28] investigated wormhole geometry in $f(Q)$ gravity. Several $f(Q)$ parameterizations have been analyzed, including observational constraints [29, 30] and investigating compact objects beyond the standard maximum mass limit [31, 32, 33, 34]. In addition, Bohmer [35], Paliathanasis [36, 37] and Khyllep et al. [38] presented a dynamical system analysis in $f(Q)$ gravity with perturbations.
@@
In cosmology, the cosmic chronometer (CC) method is utilized to determine the age and expansion rate of the Universe. The CC technique consists of three primary components: (i) defining a sample of optimal CC tracers, (ii) determining the differential age, and (iii) assessing systematic effects [39]. The Hubble parameter $H(z)$ is essential in determining the energy content of the Universe and its acceleration mechanism. The $H(z)$ estimation is mainly carried out at $z=0$. Still, there are other methods to determine $H(z)$, such as the detection of Baryonic Acoustic Oscillation (BAO) signal in the clustering of galaxies and quasars and analyzing Type Ia Supernovae (SNe Ia) observation [40, 41, 42, 43, 44]. Pantheon+ is an analysis that expands the original Pantheon framework to combine an even larger number of SNe Ia samples to understand the complete expansion history. In this study, we used the observational Hubble data (CC sample), Pantheon+, and BAO data sets to investigate the expansion history of the Universe and the behavior of other geometrical parameters.
@@
This study investigates a specific subclass of the $f(Q,B)$ model to assess its potential as an alternative to the conventional cosmological framework. We have developed a numerical approach to predict the redshift behavior of the Hubble expansion rate. Our findings indicate that while the model can replicate the low-redshift behavior of the standard $\Lambda$CDM model, it exhibits notable differences at high redshifts. The $f(Q,B)$ model emerges as a viable candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, thereby supporting its validity as an alternative explanation for the observed acceleration of the Universe. We examined the background cosmological dynamics of the selected model and evaluated its feasibility using Bayesian analysis, supported by Markov Chain Monte Carlo (MCMC) methods, applied to late-time cosmic observations, including Pantheon+, CC and BAO datasets. Additionally, we introduced a dynamical system analysis to assess the stability of the model. A significant outcome of our analysis is the identification of a stable critical point within the dynamical system using center manifold theory (CMT). This critical point corresponds to the de Sitter phase, a well-established cosmological epoch characterized by accelerated expansion. The stability of this critical point suggests that, given certain initial conditions, the Universe will inevitably move towards and remain within the de Sitter phase. This finding aligns with current observations suggesting a late-time Universe dominated by DE and undergoing accelerated expansion.
@@
In teleparallel gravity, the boundary term $B$ can be incorporated into the Lagrangian, resulting in $f(T,B)$ theories that exhibit rich phenomenology [45, 46]. However, within the framework of nonmetricity gravity, the Lagrangian of symmetric teleparallel gravity does not account for the role of $B$. This has led to the development of the $f(Q,B)$ theory, which is currently of significant interest to cosmologists [18, 47]. Our study explores the concept of an accelerating Universe by introducing a novel and straightforward parametrization for the Hubble parameter. This article is divided into five sections. Section II presents the geometrical framework of symmetric teleparallelism, also formulating $f(Q,B)$ gravity and extracting the general metric and affine connection field equations. In Section III, we apply this formulation to a cosmological setup, resulting in $f(Q,B)$ cosmology with observational data sets. Building on the model presented in Section IV, we performed a dynamical system analysis to investigate its long-term behavior and identify any stable or unstable states. Finally, Section V concludes the article with the results and discussions.
II Symmetric teleparallel gravity
@@
We examine a gravitational model defined by the four-dimensional metric tensor $g_{\mu\nu}$ and the covariant derivative ∇̊$\mathring{\nabla}_{\mu}\leavevmode\nobreak$, which is constructed using the generic connection $\Gamma_{\mu\nu}^{\zeta}\,$. Within the framework of Symmetric Teleparallel General Relativity (STGR), the connection $\Gamma_{\mu\nu}^{\zeta}$ is both flat and torsionless. Consequently, this results in $R_{;\eta\mu\nu}^{\zeta}=0$ and $\mathrm{T}_{\mu\nu}^{\eta}=0$. Furthermore, it retains the symmetries of the metric tensor $g_{\mu\nu}$. Autoparallels are defined as [48]
@@
where $\mathring{\Gamma}_{\mu\nu}^{\sigma}$, $L^{\sigma}_{\,\,\mu\nu}$, and $K^{\sigma}_{\,\,\,\mu\nu}$ respectively represent the Levi-Civita connection, the deformation tensor, and the contortion tensor. These are defined as:
@@
In symmetric teleparallel theory, we can always allowed to choose a suitable diffeomorphism that vanishes the general affine connection $\Gamma_{\mu\nu}^{\zeta}$, known as the coincident gauge [49]. As a consequence, the covariant derivative reduces to the partial derivative, and the symmetric teleparallel postulates (i.e., vanishing curvature and torsion) enforce that the general connection becomes the Levi-Civita connection, which is symmetric by construction.
@@
As a result, the nonmetricity scalar $Q$, which is defined in [50], has been introduced
@@
the traces $Q_{\mu}=Q_{\mu\nu}^{\phantom{\mu\nu}\nu}$ and $\tilde{Q}_{\mu}=Q_{\phantom{\nu}\mu\nu}^{\nu\phantom{\mu}\phantom{\mu}}$ are used in this context.
@@
The gravitational action integral for STGR is expressed as follows
∫$\int d^{4}x\sqrt{-g}Q\simeq\int d^{4}x\sqrt{-g}R-B\,,$ (12)
@@
However, the equivalence is lost when nonlinear components of the nonmetricity scalar $Q$ are introduced as in $f(Q)$–gravity in the gravitational action. Moreover, the corresponding gravitational theory has no longer dynamical equivalence with GR or its generalization, $f(\mathring{R})$–gravity.
The action integral for symmetric teleparallel $f(Q)$–gravity [49, 51] is expressed as follows
@@
II.1 $f(Q,B)$ cosmology
A recent extension of the $f(Q,B)$ theory [18, 47, 36] incorporates a boundary term into the gravitational action integral. This generalization includes the gravitational action integral in the following manner
$S=\int\left[\frac{1}{2\kappa}f(Q,B)\right]\sqrt{-g}d^{4}x,$ (14)
@@ II.1 $f(Q,B)$ cosmology
To construct a realistic cosmological model, we consider a matter action $S_{m}$, associated with the energy-momentum tensor $\Theta_{\mu\nu}$. As shown in [47], varying the total action $S+S_{m}$ leads to the following Friedmann equations:
$\kappa T_{\mu\nu}=-\frac{f}{2}g_{\mu\nu}+\frac{2}{\sqrt{-g}}\partial_{\eta}\left(\sqrt{-g}f_{Q}P^{\eta}{}_{\mu\nu}\right)\\ +(P_{\mu\alpha\beta}Q_{\nu}{}^{\alpha\beta}-2P_{\alpha\beta\nu}Q^{\alpha\beta}{}_{\mu})f_{Q}\\ +\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}-2P^{\eta}{}_{\mu\nu}\partial_{\eta}\right)f_{B}\,,$ (15)
This can be expressed in a covariant manner
$\kappa T_{\mu\nu}=-\frac{f}{2}g_{\mu\nu}+2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})+\left(\mathring{G}_{\mu\nu}+\frac{Q}{2}g_{\mu\nu}\right)f_{Q}\\ +\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\,.$ (16)
@@
$T^{\text{eff}}_{\mu\nu}=T_{\mu\nu}+\frac{1}{\kappa}\left[\frac{f}{2}g_{\mu\nu}-2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})-\frac{Qf_{Q}}{2}g_{\mu\nu}\right.\\ \left.-\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\right]\,,$ (17)
@@
In this section, we explore the application of $f(Q,B)$ gravity within a cosmological context and introduce $f(Q,B)$ cosmology. Our analysis focuses on a homogeneous and isotropic flat Friedmann-Robertson-Walker (FRW) spacetime described by the line element in cartesian coordinates
@@
Following this section, it has been demonstrated that within the context of $f(Q,B)$ gravity, an additional effective sector of geometrical origin can be obtained as shown in Eq. (17). Consequently, when considered in a cosmological context, this term can be interpreted as an effective dark-energy sector, which possesses an energy-momentum tensor
$T^{\text{DE}}_{\mu\nu}=\frac{1}{f_{Q}}\left[\frac{f}{2}g_{\mu\nu}-2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})-\frac{Qf_{Q}}{2}g_{\mu\nu}\right.$
−$\ \ \ \ \ \ \ \ \ \left.-\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\right]\,,$ (20)
@@
$3H^{2}$ $=$ $\kappa\left(\rho+\rho_{\text{DE}}\right),$ (22)
$\left(2\dot{H}+3H^{2}\right)$ $=$ −$-\kappa\left(p+p_{\text{DE}}\right),$ (23)
where $\rho$ and $p$ represent the energy density and pressure of the matter sector, respectively, treated as a perfect fluid. Additionally, we have defined the effective dark-energy density and pressure as follows:
@@
$p_{\text{DE}}=\frac{1}{\kappa}\Big{[}-2\dot{H}\left(1-f_{Q}\right)-3H^{2}\left(1-2f_{Q}\right)+\frac{f}{2}+2H\dot{f}_{Q}$
−$-\left(9H^{2}+3\dot{H}\right)f_{B}+\ddot{f_{B}}\Big{]}.$ (25)
Since standard matter is conserved independently, with $\dot{\rho}+3H\left(\rho+p\right)=0$, it can be deduced from Eqs. (24) and (II.1) that the dark energy density and pressure conform to the standard evolution equation
@@
Finally, we can define the parameter for the dark energy equation of state as
@@
II.2 Power law $f(Q,B)$
In this study, we propose a specific mathematical form of $f(Q,B)$ to capture the characteristic power-law behaviors observed in different stages of the evolution history of the Universe, i.e., at different cosmological epochs. This form is inspired by the work of Bahamonde and Capozziello [52], which utilizes the Noether Symmetry approach. The proposed form is given by:

### v5

@@ Abstract
In this work, we explore the cosmological stability of $f(Q,B)$ gravity using a dynamical system approach, where $Q$ denotes the nonmetricity scalar and $B$ represents the boundary term. We determine the model parameters of $f(Q,B)$ through Bayesian statistical analysis, employing Markov Chain Monte Carlo techniques. This analysis incorporates numerical solutions and observational data from cosmic chronometers, the extended Pantheon+ data set, and baryonic acoustic oscillation measurements. Our findings reveal a stable critical point within the dynamical system of the model, corresponding to the de Sitter phase, which is consistent with current observations of the Universe dominated by dark energy and undergoing late-time accelerated expansion. Additionally, we utilize Center Manifold Theory to examine the stability of this critical point, providing deeper insights into the behavior of the model. The cosmological implications of $f(Q,B)$ gravity indicate a smooth transition in the deceleration parameters from deceleration to the acceleration phase, underscoring the potential of the model to describe the evolution of the Universe. Our results suggests that the $f(Q,B)$ model presents a viable alternative to the standard $\Lambda$CDM model, effectively capturing the observed acceleration of the Universe and offering a robust framework for explaining the dynamics of cosmic expansion.
observations – cosmological parameters – distance scale – methods: statistical
1 Introduction
@@
The Lambda Cold Dark Matter ($\Lambda$CDM) model is widely regarded as the standard cosmological model, providing a fundamental framework for interpreting cosmic microwave background (CMB) data (Aghanim et al., 2020a). It incorporates cold dark matter, and late-time accelerated expansion driven by dark energy (DE) within a homogeneous and isotropic Universe. The cosmological constant $(\Lambda)$, integrated into Einstein’s field equations, has a robust theoretical basis. However, recent observations have highlighted discrepancies and tensions in $\Lambda$CDM model parameters across different cosmic epochs (Di Valentino et al., 2021; Perivolaropoulos & Skara, 2022; Lohakare et al., 2024), alongside persistent theoretical challenges associated with $\Lambda$ (Weinberg, 1989; Peebles & Ratra, 2003). These inconsistencies suggest that the $\Lambda$CDM model may not fully account for all missing physics. The discovery of the accelerating Universe expansion is one of the most surprising findings in cosmology, indicating either a breakdown of Einstein’s theory of gravity on cosmological scales or the dominance of DE with exotic properties. Cosmic acceleration has driven efforts to uncover its origin, with experiments measuring expansion and growth to one percent precision or better. General relativity (GR) is non-renormalizable at microscopic scales (Hooft & Veltman, 1974; Goroff et al., 1985), and the absence of a coherent quantum gravity theory necessitates modifications or extensions of GR (Hooft & Veltman, 1974; Gleiser, 2005). The standard approach develops gravity theories that encompass GR in specific scenarios while introducing additional degrees of freedom (Nojiri & Odintsov, 2011; Capozziello & De Laurentis, 2011; Bahamonde et al., 2023a; Franco et al., 2020; Heisenberg et al., 2023). These theories arise from fundamental considerations of gravitational field structures and principles (Böhmer & Jensko, 2021, 2023; Saridakis et al., 2023), with a focus on maintaining the principle of equivalence at quantum scales and addressing causal and geodesic structures. The metric requirement remains essential when considering gravity as a gauge theory.
@@
Novel classes of modified gravity theories have emerged, incorporating curvature, torsion, and nonmetricity scalars. These classes arise even though the unmodified theories are mathematically equivalent at the equation level. The key lies in the difference between the torsion scalar $T$ and the nonmetricity scalar $Q$, which deviates from the usual Levi-Civita Ricci scalar $\mathring{R}$ of GR due to additional terms: $\mathring{R}=-T+B$ and $\mathring{R}=Q+B$, respectively; where $B$ is the boundary term. All quantities denoted by $(\,\mathring{}\,)$ are computed in relation to the Levi-Civita connection $\mathring{\Gamma}$. The objects framed in GR can be identified by an over-circle symbol. A geometric trinity of gravity of second-order can be observed in $\mathring{R}$, $B-T$, $Q+B$, whereas $f(\mathring{R})$, $f(B-T)$, $f(Q+B)$ can be regarded as a geometric trinity of gravity of fourth-order (Beltrán Jiménez et al., 2019; Capozziello et al., 2023, 2022). Consequently, arbitrary functions $f(\mathring{R})$, $f(T)$, and $f(Q)$ no longer share a total derivative relationship. Furthermore, scalar fields can be introduced within this framework, leading to theories of scalar-tensor (De Felice & Tsujikawa, 2011; Horndeski, 1974), scalar torsion (Bahamonde et al., 2019, 2023b; Geng et al., 2011), and scalar nonmetricity (Rünkla & Vilson, 2018; Järv et al., 2018), each offering intriguing possibilities. Recently, Heisenberg (2023) reviewed various cosmological models in $f(Q)$ gravity. Considering energy conditions, Banerjee et al. (2021) investigated wormhole geometry in $f(Q)$ gravity. Several $f(Q)$ parameterizations have been analyzed, including observational constraints (Narawade & Mishra, 2023) and investigating compact objects beyond the standard maximum mass limit (Maurya et al., 2023, 2022a; Lohakare et al., 2023a; Maurya et al., 2022b). In addition, Böhmer et al. (2023), Paliathanasis (2024a, b) and Khyllep et al. (2023) presented a dynamical system analysis in $f(Q)$ gravity with perturbations.
@@
In cosmology, the cosmic chronometer (CC) method is utilized to determine the age and expansion rate of the Universe. The CC technique consists of three primary components: (i) defining a sample of optimal CC tracers, (ii) determining the differential age, and (iii) assessing systematic effects (Moresco et al., 2022). The Hubble parameter $H(z)$ is essential in determining the energy content of the Universe and its acceleration mechanism. The $H(z)$ estimation is mainly carried out at $z=0$. Still, there are other methods to determine $H(z)$, such as the detection of baryonic acoustic oscillation (BAO) signal in the clustering of galaxies and quasars and analyzing Type Ia Supernovae (SNe Ia) observation (Riess et al., 2021; Font-Ribera et al., 2014; Raichoor et al., 2020; Hou et al., 2020; Riess et al., 2018). Pantheon+ is an analysis that expands the original Pantheon framework to combine an even larger number of SNe Ia samples to understand the complete expansion history. In this study, we used the observational Hubble data (CC sample), Pantheon+, and BAO data sets to investigate the expansion history of the Universe and the behavior of other geometrical parameters.
@@
This study investigates a specific subclass of the $f(Q,B)$ model to assess its potential as an alternative to the conventional cosmological framework. We have developed a numerical approach to predict the redshift behavior of the Hubble expansion rate. Our findings indicate that while the model can replicate the low-redshift behavior of the standard $\Lambda$CDM model, it exhibits notable differences at high redshifts. The $f(Q,B)$ model emerges as a viable candidate for explaining the current epochs and effectively captures the evolution of energy components over cosmic time, thereby supporting its validity as an alternative explanation for the observed acceleration of the Universe. We examined the background cosmological dynamics of the selected model and evaluated its feasibility using Bayesian analysis, supported by Markov Chain Monte Carlo (MCMC) methods, applied to late-time cosmic observations, including the Pantheon+, CC and BAO data sets. Additionally, we introduced a dynamical system analysis to assess the stability of the model. A significant outcome of our analysis is the identification of a stable critical point within the dynamical system using center manifold theory (CMT). This critical point corresponds to the de Sitter phase, a well-established cosmological epoch characterized by accelerated expansion. The stability of this critical point suggests that, given certain initial conditions, the Universe will inevitably move toward and remain within the de Sitter phase. This finding aligns with current observations suggesting a late-time Universe dominated by DE and undergoing accelerated expansion.
@@
In teleparallel gravity, the boundary term $B$ can be incorporated into the Lagrangian, resulting in $f(T,B)$ theories that exhibit rich phenomenology (Bahamonde et al., 2015; Kadam et al., 2023). However, within the framework of nonmetricity gravity, the Lagrangian of symmetric teleparallel gravity does not account for the role of $B$. This has led to the development of the $f(Q,B)$ theory, which is currently of significant interest to cosmologists (Capozziello et al., 2023; De et al., 2024). Our study explores the concept of an accelerating Universe by introducing a novel and straightforward parametrization for the Hubble parameter. This article is divided into five sections. Section 2 presents the geometrical framework of symmetric teleparallelism, also formulating $f(Q,B)$ gravity and extracting the general metric and affine connection field equations. In Section 3, we apply this formulation to a cosmological setup, resulting in $f(Q,B)$ cosmology with observational data sets. Building on the model presented in Section 4, we perform a dynamical system analysis to investigate its long-term behavior and identify any stable or unstable states. Finally, Section 5 concludes the article with the results and discussions.
2 Symmetric teleparallel gravity
@@
We examine a gravitational model defined by the four-dimensional metric tensor $g_{\mu\nu}$ and the covariant derivative ̊$\mathring{\nabla}_{\mu}~{}$, which is constructed using the generic connection $\Gamma_{\mu\nu}^{\zeta}\,$. Within the framework of Symmetric Teleparallel General Relativity (STGR), the connection $\Gamma_{\mu\nu}^{\zeta}$ is both flat and torsionless. Consequently, this results in $R_{;\eta\mu\nu}^{\zeta}=0$ and $\mathrm{T}_{\mu\nu}^{\eta}=0$. Furthermore, it retains the symmetries of the metric tensor $g_{\mu\nu}$. Autoparallels are defined as (Obukhov & Puetzfeld, 2021)
@@
where $\mathring{\Gamma}_{\mu\nu}^{\sigma}$, $L^{\sigma}_{\,\,\mu\nu}$, and $K^{\sigma}_{\,\,\,\mu\nu}$ respectively represent the Levi-Civita connection, the deformation tensor, and the contortion tensor. These are defined as
@@
In symmetric teleparallel theory, we are always allowed to choose a suitable diffeomorphism that vanishes the general affine connection $\Gamma_{\mu\nu}^{\zeta}$, known as the coincident gauge (Jiménez et al., 2018). As a consequence, the covariant derivative reduces to the partial derivative, and the symmetric teleparallel postulates (i.e., vanishing curvature and torsion) enforce that the general connection becomes the Levi-Civita connection, which is symmetric by construction.
@@
As a result, the nonmetricity scalar $Q$, which is defined in (Nester & Yo, 1999), has been introduced:
@@
and the traces $Q_{\mu}=Q_{\mu\nu}^{\phantom{\mu\nu}\nu}$ and $\tilde{Q}_{\mu}=Q_{\phantom{\nu}\mu\nu}^{\nu\phantom{\mu}\phantom{\mu}}$ are used in this context.
@@
The gravitational action integral for STGR is expressed as follows:
$\int d^{4}x\sqrt{-g}Q\simeq\int d^{4}x\sqrt{-g}R-B\,,$ (12)
@@
However, the equivalence is lost when nonlinear components of the nonmetricity scalar $Q$ are introduced as in $f(Q)$–gravity in the gravitational action. Moreover, the corresponding gravitational theory no longer has dynamical equivalence with GR or its generalization, $f(\mathring{R})$ gravity.
The action integral for symmetric teleparallel $f(Q)$ gravity (Jiménez et al., 2018, 2020) is expressed as follows:
@@
2.1 $f(Q,B)$ cosmology
A recent extension of the $f(Q,B)$ theory (Capozziello et al., 2023; De et al., 2024; Paliathanasis, 2024a) incorporates a boundary term into the gravitational action integral. This generalization includes the gravitational action integral in the following manner:
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2\kappa}f(Q,B)+\mathcal{L}_{m}\right],$ (14)
@@ 2.1 $f(Q,B)$ cosmology
To construct a realistic cosmological model, we consider a matter action $S_{m}$, associated with the energy-momentum tensor $\Theta_{\mu\nu}$. As shown in (De et al., 2024), varying the total action $S$ leads to the following Friedmann equations:
$\kappa T_{\mu\nu}=-\frac{f}{2}g_{\mu\nu}+\frac{2}{\sqrt{-g}}\partial_{\eta}\left(\sqrt{-g}f_{Q}P^{\eta}{}_{\mu\nu}\right)+(P_{\mu\alpha\beta}Q_{\nu}{}^{\alpha\beta}-2P_{\alpha\beta\nu}Q^{\alpha\beta}{}_{\mu})f_{Q}+\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}-2P^{\eta}{}_{\mu\nu}\partial_{\eta}\right)f_{B}\,,$
This can be expressed in a covariant manner:
$\kappa T_{\mu\nu}=-\frac{f}{2}g_{\mu\nu}+2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})+\left(\mathring{G}_{\mu\nu}+\frac{Q}{2}g_{\mu\nu}\right)f_{Q}+\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\,.$ (16)
@@
$T^{\text{eff}}_{\mu\nu}=T_{\mu\nu}+\frac{1}{\kappa}\left[\frac{f}{2}g_{\mu\nu}-2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})-\frac{Qf_{Q}}{2}g_{\mu\nu}-\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\right]\,.$ (17)
@@
In this section, we explore the application of $f(Q,B)$ gravity within a cosmological context and introduce $f(Q,B)$ cosmology. Our analysis focuses on a homogeneous and isotropic flat Friedmann–Lemaítre–Robertson–Walker (FLRW) spacetime described by the line element in Cartesian coordinates
@@
Following this section, it has been demonstrated that within the context of $f(Q,B)$ gravity, an additional effective sector of geometrical origin can be obtained as shown in Equation (17). Consequently, when considered in a cosmological context, this term can be interpreted as an effective DE sector, which possesses an energy-momentum tensor
$T^{\text{DE}}_{\mu\nu}=\frac{1}{f_{Q}}\left[\frac{f}{2}g_{\mu\nu}-2P^{\eta}{}_{\mu\nu}\mathring{\nabla}_{\eta}(f_{Q}-f_{B})-\frac{Qf_{Q}}{2}g_{\mu\nu}-\left(\frac{B}{2}g_{\mu\nu}-\mathring{\nabla}_{\mu}\mathring{\nabla}_{\nu}+g_{\mu\nu}\mathring{\nabla}^{\alpha}\mathring{\nabla}_{\alpha}\right)f_{B}\right]\,,$ (20)
@@
$3H^{2}$ $=$ $\kappa\left(\rho_{\text{m}}+\rho_{\text{r}}+\rho_{\text{DE}}\right)\,,$ (22)
$-2\dot{H}-3H^{2}$ $=$ $\kappa\left(\frac{\rho_{\text{r}}}{3}+p_{\text{DE}}\right)\,,$ (23)
where $\rho_{\text{m}}$, $\rho_{\text{r}}$, $\rho_{\text{DE}}$ and $p_{\text{DE}}$ denote the matter density, radiation density, DE density, and DE pressure, respectively, treated as a perfect fluid. In the absence of interactions between nonrelativistic matter and radiation, each component independently adheres to its respective conservation laws, expressed as $\dot{\rho}_{\text{m}}+3H\rho_{\text{m}}=0$ for matter and $\dot{\rho}_{\text{r}}+4H\rho_{\text{r}}=0$ for radiation. Additionally, we have defined the effective DE density and pressure as follows:
@@
$p_{\text{DE}}=\frac{1}{\kappa}\Big{[}-2\dot{H}\left(1-f_{Q}\right)-3H^{2}\left(1-2f_{Q}\right)+\frac{f}{2}+2H\dot{f}_{Q}-\left(9H^{2}+3\dot{H}\right)f_{B}+\ddot{f_{B}}\Big{]}.$ (25)
Since standard matter is conserved independently, it can be deduced from Equations (24) and (25) that the DE density and pressure conform to the standard evolution equation
@@
Finally, we can define the parameter for the DE equation of state (EoS) as
@@
2.2 Power law $f(Q,B)$
In this study, we propose a specific mathematical form of $f(Q,B)$ to capture the characteristic power-law behaviors observed in different stages of the evolution history of the Universe, i.e., at different cosmological epochs. This form is inspired by the work of Bahamonde & Capozziello (2017), which utilizes the Noether symmetry approach. The proposed form is given by: